OpenStudy (anonymous):

Double integral (x+y)dxdy bounded by x+y=2 and y=x^2. I know the imits for x(sqrt(y) to y-2), but I am not sure about limits of y.

OpenStudy (anonymous):

?any help

OpenStudy (nikvist):

\[-2\leq x\leq 1\quad,\quad x^2\leq y\leq 2-x\]

OpenStudy (anonymous):

yh but what you have is dydx. I want dxdxy

OpenStudy (nikvist):

this is easier way

OpenStudy (anonymous):

i know it is but my professor wants it dxdy

OpenStudy (anonymous):

i think y is 1 to 4 may be

OpenStudy (nikvist):

\[0\leq y\leq 1\quad,\quad -\sqrt{y}\leq x\leq\sqrt{y}\]\[1\leq y\leq 4\quad,-\sqrt{y}\leq x\leq 2-y\]

OpenStudy (anonymous):

thank you:) This works